a curve is such that dy/dx = -x square + 5x - 4.
1) find the x- coordinates of each of the stationary points of the curve.
2) obtain an expression for d square y/d x square (second derivative) and hence or otherwise find the nature of each of the stationary points .
3)given that the curve passes through the point (6 , 2 ), find the equation of the curve .
1
Expert's answer
2020-04-22T18:43:54-0400

1) To find coordinates of stationary points we should solve "\frac{dy}{dx}=0"

"-x^2+5x-4=0"

"-(x-1)(x-4)=0"

"x=1" or "x=4"


2) "\frac{d^2y}{dx^2}=(-x^2+5x-4)'=-2x+5"

when "x=1" "\frac{d^2y}{dx^2}=3>0" point of local minimum

when "x=4" "\frac{d^2y}{dx^2}=-3<0" point of local maximum


3) "y=\int(-x^2+5x-4)dx=\frac{-x^3}{3}+5\frac{x^2}{2}-4x+C"

"(6,2)"

"2=-\frac{6^3}{3}+5*\frac{6^2}{2}-4*6+C"

"2= -72+90-24+C"

"C=8"

"y=\frac{-x^3}{3}+5\frac{x^2}{2}-4x+8"